The ages of a certain group of students in a junior high school follow a normal

distribution with a mean of 13 and a standard deviation of 1.5. Find the probability

that a randomly selected student is between 10-16 years old.


1
Expert's answer
2022-04-06T09:20:59-0400

mean =13

standard deviation=1.5

P(10<x<16)=P(z2"\le"16)-P(z1"\le" 10)

z="\frac{x-\mu}{\sigma }"


z2= "\frac{16-13}{1.5}"

z2=2

P(z2"\le" 2)=0.9772


z1="\frac{10-13}{1.5}" =-2

P(z1"\le" -2)=0.0228

Thus

P(10<x<16)=0.9772-0.0228

P(10<x<16)=0.9544

The probability that a randomly selected student is between 10-16 years old is 0.9544.